The perimeter of a rectangle is 62 inches. If the length of the rectangle is 1 inch more than twice the width what are the dimensions of the rectangle?
I need to set up an equation and solve

Answer by rfer(16322)

2x+2(2x+1)=62
2x+4x+2=62
6x=60
x=10
2x+1=21